# Talking to DINO: Bridging Self-Supervised Vision Backbones with Language for Open-Vocabulary Segmentation

### v1

@@ 4.1 Experimental Setup
Datasets. We evaluate our approach on eight widely-used semantic segmentation benchmarks, which we categorize based on the inclusion of a background class. Specifically, we conduct experiments on the validation sets of Pascal VOC 2012 [15], Pascal Context [30], COCO Stuff [7], Cityscapes [11], and ADE20K [56, 57], that contain 20, 59, 171, 150, and 19 semantic categories respectively and do not include the “background” class. We report additional experiments on the COCO Objects dataset [7], which consists of 80 different foreground object classes, and on modified versions of Pascal VOC 2012 and Pascal Context in which the “background” category is, instead, included.
Implementation Details. For the main experiments, we employ DINOv2 ViT-B/14 as the base model and DINOv2 ViT-L/14 as the large model, both with the CLIP ViT-B/16 text encoder. We use the DINOv2 variant with registers [12] since our method benefits from the removal of artifacts in self-attention maps [^1] A detailed analysis on the impact of using DINOv2 with registers is reported in the supplementary material.. We resize the input image to $518\times 518$ to match the original training resolution of DINOv2, obtaining $37\times 37$ patches. We train the model with the Adam optimizer, a batch size of $128$, and a learning rate of $1\times 10^{-4}$ for a total of 100 epochs on the COCO Captions 2014 training split [27], composed of around 80k images.
@@
Evaluation Protocol. We follow the standard evaluation protocol for unsupervised OVS [9], where prior access to the target data before evaluation is not allowed, and use the default class names provided by the MMSegmentation toolbox. All models are evaluated using mean Intersection-over-Union (mIoU) on all the classes of each dataset. The images are resized to have a shorter side of 448, using a sliding window approach with a stride of 224 pixels.
@@ 4.2 Comparison with the State of the Art
We compare Talk2DINO with previous state-of-the-art approaches for unsupervised OVS on the five benchmarks that do not include the “background” category and the three benchmarks with the “background” category. We consider as competitors: (i) prototype-based approaches, such as ReCo [37], OVDiff [22], FOSSIL [2], and FreeDA [3], which aim to create visual prototypes associated with the textual categories, (ii) CLIP adaptations, as MaskCLIP [58], CLIP-DIY [46], SCLIP [42], and NACLIP [19], which propose architectural modifications to enhance the localization properties of CLIP, (iii) methods trained on large-scale sets of image-captions pairs with loss functions designed to force the segmentation capabilities to emerge, like GroupViT [49], ViewCo [36], TCL [9], OVSegmentor [50], SegCLIP [28], and SILC [32], and (iv) methods that aim to combine the complementary properties of CLIP and DINO, like CLIP-DINOiser [47] and ProxyCLIP [23].
Table 1 reports the results on the five benchmarks without background (i.e., Pascal VOC, Pascal Context, COCOStuff, Cityscapes, and ADE). Specifically, we report the performance of both the base and large configurations of both Talk2DINO and the competitors, according to their definitions in the original papers. Moreover, we divide the table into two sections depending on whether a mask refinement technique is employed. Specifically, FreeDA exploits Felzenszwalb’s superpixel algorithm [17], FOSSIL introduces a custom region proposal algorithm named OpenCut, while all the other methods refine their masks with PAMR. As it can be observed, our approach achieves the best average mIoU on all the configurations and presents a consistent improvement compared to the considered competitors, with and without the mask refinement, across all datasets.
@@
Model Visual Backbone Frozen VOC Context Object
ViT-Base
GroupViT [49] Custom ViT 51.1 19.0 27.9
ViewCo [36] Custom ViT 52.4 23.0 23.5
OVSegmentor [50] DINO 53.8 20.4 25.1
SegCLIP [28] CLIP 52.6 24.7 26.5
TCL [9] CLIP 55.0 30.4 31.6
MaskCLIP [58] CLIP 37.2 22.6 18.9
CLIP-DIY [46] CLIP 59.9 19.7 31.0
OVDiff [22] CLIP+DINO+SD 68.4 31.2 36.2
SCLIP [42] CLIP 61.7 31.5 32.1
CLIP-DINOiser [47] CLIP 62.1 32.4 34.8
NACLIP [19] CLIP 64.1 35.0 36.2
FreeDA [3] DINOv2 52.9 35.5 25.9
FreeDA [3] CLIP+DINOv2 55.4 37.9 36.2
ProxyCLIP [23] CLIP+DINOv2 58.6 33.8 37.4
ProxyCLIP [23] CLIP+DINO 61.3 35.3 37.5
Talk2DINO (Ours) DINOv2 65.9 38.4 44.6
ViT-Large
NACLIP [19] CLIP 57.9 36.4 34.6
FreeDA [3] DINOv2 50.3 34.3 26.7
FreeDA [3] CLIP+DINOv2 55.4 38.3 37.4
ProxyCLIP [23] CLIP+DINOv2 56.6 33.0 36.7
ProxyCLIP [23] CLIP+DINO 60.6 34.5 39.2
Talk2DINO (Ours) DINOv2 66.6 37.7 43.0
Table 2: Comparison with unsupervised OVS models on Pascal VOC [15], Pascal Context [30], and COCO Object [7], considering the background category. For each method, the visual backbone used is specified, along with whether it is frozen or fine-tuned. We report all methods in their best configuration.
Table 2, instead, shows the results on the three benchmarks that require assigning the “background” category to pixels that do not belong to the set of provided categories. Pascal VOC and COCO Objects present only foreground categories, also referred to as “things” in the literature, while Pascal Context presents both categories from foreground and background, also mentioned as “stuff”. Hence, following CLIP-DINOiser [47], we report the performance with the background cleaning procedure only on Pascal VOC and COCO Objects. From the results, it can be seen that Talk2DINO achieves the best or second-best result on all considered configurations. We also notice that OVDiff builds specific negative prototypes for each category to enhance its representation of the background.
The most straightforward comparison is the one with FreeDA without global similarity (i.e., with DINOv2 only as visual backbone). Indeed, this approach builds a bridge between DINOv2 and the CLIP text encoder by retrieving from a collection of visual and textual embedding pairs and by building visual prototypes for each textual category. The significant improvement achieved by Talk2DINO demonstrates that training a direct projection from the CLIP text encoder to DINOv2 leads to a more accurate bridge between the two embedding spaces without the overhead in computation and memory provided by the retrieval procedure.
@@ 4.3 Ablation Studies and Analyses
Choosing Different Visual Backbones. In Table 3 we show the performance of our approach when varying the visual backbone and the size of the employed ViT architecture. In these experiments, we keep the original resolution of the models during the training procedure and follow the same downsampling protocol employed for Talk2DINO at test time. We observe that backbones that differ from DINOv2 present unsatisfactory results and can not be aligned to the CLIP textual encoder with a learnable mapping. In particular, while DINO achieves the second best performance on average, Talk2DINO heavily benefits from the strong semantic representation of the dense features of DINOv2 and on the capabilities of its self-attention heads in highlighting coherent regions of the image – properties which are not reflected in other visual backbones (see supplementary materials for detailed analysis). Additionally, we notice that our approach demonstrates consistent and high performance across different ViT sizes, with strong results even when using the ViT-Small backbone.
@@
Visual Backbone VOC Context Stuff Cityscapes ADE
DINO ViT-S 25.8 12.6 7.0 15.2 4.9
DINOv2 ViT-S 86.8 35.3 24.6 28.3 17.0
MAE ViT-B 5.3 2.7 1.4 4.7 0.9
CLIP ViT-B 47.3 9.7 9.8 3.3 2.4
DINO ViT-B 22.0 12.3 6.9 16.4 4.7
DINOv2 ViT-B 85.3 40.5 27.9 38.2 21.8
MAE ViT-L 6.7 2.7 1.4 4.6 0.9
CLIP ViT-L 16.3 4.9 7.4 1.0 2.0
DINOv2 ViT-L 84.7 39.4 27.2 36.3 21.2
mIoU
VOC Context Stuff Cityscapes ADE
@@
Single FC (text only) 84.5 38.5 26.9 35.3 20.1
Our mapping (both vision and text) 56.8 23.9 16.8 21.9 12.6
Our mapping (vision only) 77.1 30.6 20.9 22.5 15.4
Our mapping (text only) 85.3 40.5 27.9 38.2 21.8
@@
CLS only (without self-attention) 85.4 30.4 22.5 22.9 16.4
Standard average 89.2 38.6 26.0 34.8 20.4
CLS-weighted average 87.4 34.9 22.8 29.5 16.9
CLS similarity-weighted sampling 88.2 32.4 22.7 29.9 18.8
Max CLS similarity 85.3 40.5 27.9 38.2 21.8
@@
Impact of the Proposed Components. Table 4 reports the results of Talk2DINO evaluating the impact of its core components on the overall performance. Specifically, in the first section of the table, we analyze the effect of the adopted projection $\psi$. Noticeably, replacing it with a linear projection leads to a slight drop in performance. The good performance obtained by a linear transform evidences how the DINOv2 and CLIP spaces are intrinsically compatible, as the former can be obtained through an affine transform of the latter without losing too much information. Interestingly, applying the proposed projection on top of DINOv2 or applying two projections on both spaces leads to significantly lower performance, confirming the appropriateness of the proposed approach.
In the second section of the table, we instead study the effect of the selection and aggregation strategy of self-attention heads $A_{i},i=1,...,N$ during training. In particular, we test aligning (i) directly the visual CLS token to the textual CLS token, (ii) the visual embedding from the standard average self-attention, (iii) the weighted mean of the head embeddings $v_{A_{i}}$, where the weights are given by their softmaxed similarity with the textual CLS token, (iv) a strategy in which we sample a single head embedding, where the sampling probability is given by the softmaxed similarity with the CLS token, and (v) our adopted solution in which we select the head embedding which is the most similar to the textual CLS token. The results show that only on the Pascal VOC dataset the embedding from the standard average self-attention presents improved performance. On all other benchmarks, our approach proves most effective, further validating the robustness of our selection method.
@@
Mask Refinement VOC Context Object
without background cleaning 57.3 35.7 36.5
with background cleaning 61.1 35.3 40.4
without background cleaning 61.2 38.4 39.7
with background cleaning 65.9 37.6 44.6
Table 5: Ablation study comparing results with and without the background cleaning procedure. We report the results using the base model of DINOv2.
Effect of Background Cleaning. Table 5 shows how the performance is affected by the background cleaning mechanism and by the usage of PAMR for mask refinement. It can be observed that, according to Sec. 4.2, background cleaning has a positive impact on Pascal VOC and COCO Object, which contain only foreground categories, while slightly degrades the performance on Pascal Context, which however contains categories that are usually part of the background of an image (e.g., sky, road, and wall). Further, it can be noticed that the effectiveness of the proposed background cleaning procedure is confirmed both with and without mask refinement.
Qualitative Results. Fig. 4 depicts qualitative segmentation results, in which we highlight the segmentation capabilities of Talk2DINO along with other state-of-the-art models [3, 23, 47]. We show two images from Pascal VOC and COCO Object, in which it can also be appreciated how the background cleaning procedure leads to high-quality masks and localization. Moreover, we show two images from Cityscapes and Pascal Context where Talk2DINO effectively segments small objects in the scene, such as the traffic signs, and “stuff” categories, such as sky and road.
@@ 5 Conclusion
In this paper we introduced Talk2DINO, a novel approach for open-vocabulary segmentation that bridges the spatially detailed embeddings of the DINOv2 self-supervised vision backbone with the highly semantic CLIP’s text embeddings. Our method achieves fine-grained alignment between textual concepts and visual patches without the need for extensive fine-tuning of the backbone networks, only leveraging DINOv2’s self-attention maps and a lightweight language-to-vision mapping layer. Talk2DINO achieves state-of-the-art results in standard OVS benchmarks, outperforming previous prototype-based, CLIP adaptation, and contrastive learning approaches. Our study highlights the potential of integrating vision-only and multimodal models, suggesting broader applications in fine-grained and cross-modal tasks.
@@ Acknowledgments
Figure 5: Performance vs. Parameter Count. The y-axis denotes the obtained mIoU averaged over all the five benchmarks reported in the main paper. Dashed lines connect methods tested with multiple backbone sizes; these are labeled S/B/L for Small/Base/Large ViT models. Talk2DINO offers the best trade-off.
@@ Appendix A Additional Experiments and Analyses
Analysis of Model Parameters. Fig. 5 reports a comparison of the relationship between the performance, in terms of average mIoU, and the number of parameters of the models. As it can be observed, Talk2DINO presents a lower number of parameters than the recent competitors FreeDA [3] and ProxyCLIP [23], along with an improved average mIoU. Models with a comparable number of parameters, such as TCL [9], GroupViT [49], and MaskCLIP [58], exhibit a lower performance compared to Talk2DINO. Finally, it shall be noted that models such as FreeDA, ReCo [37], and OVDiff [22] require maintaining external sources of knowledge, which increases memory consumption.
Role of DINO Registers. The main configuration of Talk2DINO, with both the base and large sizes, leverages the variant of DINOv2 with registers. Registers are a recently proposed mechanism (see Darcet et al. [12]) to mitigate the presence of artifacts in the feature maps of ViT-based backbones [14]. Artifacts are tokens that exhibit a significantly higher norm with respect to the other tokens and retain less information about their original position in the image. This phenomenon has been especially observed after long training and is more pronounced in backbones with larger sizes. In Fig. 8 we depict, on the first row, the average self-attentions between the CLS and the other tokens for the ViT-S, ViT-B, and ViT-L architectures with and without registers, while in the following rows, we show the various self-attention heads for each backbone. It can be observed that in the ViT-S the artifacts are not present, and the average self-attention between the model with and without the registers is nearly identical. Instead, the ViT-B exhibits artifacts in the top left corner, resulting in an average self-attention that is especially focused on that portion and less active on the people. This side effect is even more noticeable with the ViT-L, for which the artifact is the only visible token in the average self-attention.
In Table 6, we compare the performance achieved when employing the DINOv2 visual backbone with and without registers. The results show a downgrade in performance without the registers that is directly related to the presence of the artifacts in the self-attentions, according to the observations from Fig. 8. Indeed, the largest difference in performance is measured in the ViT-L architecture, while in the ViT-S case, the backbone without registers performs better on four benchmarks out of five. The phenomenon of the artifacts represents a threat to our approach because it limits the mechanism of selection of the self-attention head most similar to the caption, forcing the alignment between the unwanted artifact tokens and the caption.
@@
Choosing Different Visual Backbones. In Table 3 of the main paper, we reported the performance of our approach applied to different visual backbones. The results demonstrated that our training pipeline is especially suitable for DINOv2, while it leads to unsatisfactory performance on DINO, MAE, and CLIP. In this paragraph, we further explore, from a qualitative perspective, the limitations of these probed visual backbones. Fig. 9 showcases the average self-attentions between the CLS and the other tokens on the first row, breaking down the contributions from the various self-attention heads in the successive rows. We can notice that the self-attention heads of CLIP present a noise that, similar to what we observed for DINOv2 without registers, represents a limitation for our proposed training procedure. On the contrary, the self-attention heads of DINO and MAE are clean and highlight homogeneous regions of the image. In these cases, the performance gap with DINOv2 can be related to the insufficient semantic richness of the dense features of these backbones. These insights are also confirmed by the ablation study reported in FreeDA [3], in which a similar performance gap has been observed when using the CLIP and DINO backbones instead of DINOv2.
@@
Visual Backbone VOC Context Stuff Cityscapes ADE
DINOv2 (without registers) ViT-S 83.7 38.3 25.7 32.9 19.8
DINOv2 (with registers) ViT-S 86.8 35.3 24.6 28.3 17.0
DINOv2 (without registers) ViT-B 68.3 28.5 20.5 20.9 13.6
DINOv2 (with registers) ViT-B 85.3 40.5 27.9 38.2 21.8
DINOv2 (without registers) ViT-L 58.6 21.2 15.8 14.5 8.9
DINOv2 (with registers) ViT-L 84.7 39.4 27.2 36.3 21.2
Table 6: Performance comparison of DINOv2 backbones with and without registers.

### v3

@@ 4.1 Experimental Setup
Datasets. We evaluate our approach on eight widely-used semantic segmentation benchmarks, which we categorize based on the inclusion of a background class. Specifically, we conduct experiments on the validation sets of Pascal VOC 2012 [15], Pascal Context [34], COCO Stuff [7], Cityscapes [11], and ADE20K [61, 62], that contain 20, 59, 171, 150, and 19 semantic categories respectively and do not include the “background” class. We report additional experiments on the COCO Objects dataset [7], which consists of 80 different foreground object classes, and on modified versions of Pascal VOC 2012 and Pascal Context in which the “background” category is, instead, included (i.e., with 21 and 60 semantic categories respectively).
Implementation Details. For the main experiments, we employ DINOv2 ViT-B/14 as the base model and DINOv2 ViT-L/14 as the large model, both with the CLIP ViT-B/16 text encoder. We use the DINOv2 variant with registers [12] since our method benefits from the removal of artifacts in self-attention maps. We train the model with the Adam optimizer, a batch size of $128$, and a learning rate of $1\times 10^{-4}$ for a total of 100 epochs on the COCO Captions 2014 training split [31], composed of around 80k images.
@@
Evaluation Protocol. We follow the standard evaluation protocol for unsupervised OVS [9], where prior access to the target data before evaluation is not allowed, and use the default class names provided by the MMSegmentation toolbox. The images are resized to have a shorter side of 448, using a sliding window approach with a stride of 224 pixels. All models are evaluated using mean Intersection-over-Union (mIoU) on all the classes of each dataset.
@@ 4.2 Comparison with the State of the Art
We compare Talk2DINO with previous state-of-the-art approaches for unsupervised OVS on the five benchmarks that do not include the “background” category and the three benchmarks with the “background” category. We consider as competitors: (i) prototype-based approaches, such as ReCo [41] and FreeDA [3], which aim to create visual prototypes associated with the textual categories, (ii) CLIP adaptations, as MaskCLIP [63], CLIP-DIY [51], SCLIP [47], ClearCLIP [25], and NACLIP [18], which propose architectural modifications to enhance its localization properties, (iii) methods trained on sets of image-caption pairs with objectives designed to force the segmentation capabilities to emerge, like GroupViT [54], TCL [9], SILC [36], and dino.txt [21], and (iv) methods that aim to combine the properties of CLIP and DINO, as CLIP-DINOiser [52], LaVG [22], and ProxyCLIP [26].
Table 1 reports the results on the five benchmarks without background (i.e., Pascal VOC-20, Pascal Context-59, COCO Stuff, Cityscapes, and ADE) and the three benchmarks with background (i.e., Pascal VOC-21, Pascal Context-60, and COCO Object). Specifically, we report the performance of both the base and large configurations of both Talk2DINO and the competitors, according to their definitions in the original papers. Moreover, we divide the table into two sections depending on whether a mask refinement technique is employed. Specifically, LaVG exploits a custom region proposer combined with DenseCRF [24], while all the other methods refine their masks with PAMR. Regarding datasets with background, Pascal VOC and COCO Objects present only foreground categories, also referred to as “things” in the literature, while Pascal Context presents both categories from foreground and background, also mentioned as “stuff”. Hence, following CLIP-DINOiser [52], we report the performance with the background cleaning procedure described in Sec. 3.3 only on Pascal VOC and COCO Objects.
As it can be observed, our approach achieves the best average mIoU on all the configurations and presents a consistent improvement compared to the considered competitors, with and without the mask refinement, across all datasets except Cityscapes. The most straightforward comparison is the one with FreeDA without global similarity (i.e., with DINOv2 only as visual backbone). It builds a bridge between DINOv2 and the CLIP text encoder by retrieving from a collection of visual-textual embedding pairs and by building visual prototypes for each textual category. The significant improvement achieved by Talk2DINO demonstrates that training a direct projection from the CLIP text encoder to DINOv2 leads to a more accurate bridge between the two embedding spaces without the overhead in computation and memory provided by the retrieval procedure.
@@
Visual Backbone V20 C59 Stuff City ADE
DINO ViT-S 27.7 13.2 7.4 14.8 5.2
DINOv2 (without registers) ViT-S 83.7 38.3 25.8 32.9 19.8
DINOv2 (with registers) ViT-S 86.9 35.3 24.5 27.2 16.9
MAE ViT-B 9.5 4.3 2.0 4.0 1.1
CLIP ViT-B 55.4 14.9 12.2 6.2 3.8
DINO ViT-B 27.3 11.2 7.9 13.6 4.5
DINOv2 (without registers) ViT-B 74.2 31.9 23.0 27.9 16.5
DINOv2 (with registers) ViT-B 87.1 39.8 28.1 36.6 21.1
MAE ViT-L 6.7 2.7 1.4 4.6 0.9
CLIP ViT-L 16.6 5.0 7.7 0.9 2.0
DINOv2 (without registers) ViT-L 56.0 20.1 14.9 18.1 8.2
DINOv2 (with registers) ViT-L 87.1 39.1 27.0 35.8 21.1
@@ 4.3 Ablation Studies and Analyses
Choosing Different Visual Backbones. In Table 2 we show the performance of our approach when varying the visual backbone and the size of the employed ViT architecture. We observe that backbones that differ from DINOv2 present unsatisfactory results and can not be aligned to the CLIP textual encoder with a learnable mapping. In particular, while DINO achieves the second best performance on average, Talk2DINO heavily benefits from the strong semantic representation of the dense features of DINOv2 and on the capabilities of its self-attention heads in highlighting coherent regions of the image – properties which are not reflected in other visual backbones. We refer to the supplementary materials for more details on self-attention heads.
Moreover, our results emphasize the critical role of registers [12] in DINOv2, as demonstrated by the comparison between its variants with and without registers. Registers are a recently proposed mechanism to mitigate the presence of artifacts in the feature maps of ViT-based backbones. Artifacts are tokens that exhibit a significantly higher norm with respect to the other tokens and retain less information about their original position in the image. The alignment process in our method relies on high-quality attention maps, and the presence of artifacts poses a challenge by limiting the selection of the most relevant self-attention heads. Interestingly, since register-related artifacts are more pronounced in larger backbones, the ViT-S variant without registers maintains competitive performance compared to its register-enabled counterpart. Finally, we observe that our approach maintains robust and consistent performance across different ViT sizes, achieving strong results even with the compact ViT-S backbone. This suggests that our method is effective across a range of model sizes, making it adaptable to varying computational constraints. Additional insights and detailed evaluations of the different visual backbones can be found in the supplementary.
@@
V20 C59 Stuff City ADE
@@
Linear Projection (text only) 85.1 37.9 26.7 35.6 20.1
Our mapping (both vision and text) 59.2 27.3 18.9 23.5 13.5
Our mapping (vision only) 84.6 35.2 26.2 20.4 15.5
Our mapping (text only) 87.1 39.8 28.1 36.6 21.1
@@
CLS only (without self-attention) 84.5 30.6 23.0 22.6 17.2
Standard average 89.6 36.9 25.6 33.5 19.7
CLS-weighted average 87.6 35.2 23.1 29.3 17.5
CLS similarity-weighted sampling 88.2 32.9 22.6 27.0 17.9
Max CLS similarity 87.1 39.8 28.1 36.6 21.1
@@
Impact of the Proposed Components. Table 3 reports the results of Talk2DINO evaluating the impact of its core components on the overall performance. Specifically, in the first section of the table, we analyze the effect of the adopted projection $\psi$. Replacing it with a linear projection leads to a slight performance drop. The good performance obtained by a linear transform evidences how the DINOv2 and CLIP spaces are intrinsically compatible, as the former can be obtained through an affine transform of the latter without losing too much information. Interestingly, applying the proposed projection on top of DINOv2 or using two projections on both spaces significantly lowers performance, confirming the appropriateness of the proposed approach.
In the second section of the table, we instead study the effect of the selection and aggregation strategy of self-attention heads $A_{i},i=1,...,N$ during training. In particular, we test aligning (i) directly the visual CLS token to the textual CLS token, (ii) the visual embedding from the standard average self-attention, (iii) the weighted mean of the head embeddings $v_{A_{i}}$, where the weights are given by their softmaxed similarity with the textual CLS token, (iv) a strategy in which we sample a single head embedding, where the sampling probability is given by the softmaxed similarity with the CLS token, and (v) our adopted solution in which we select the head embedding which is the most similar to the textual CLS token. The results show that only on the Pascal VOC dataset – composed mostly by large subjects in foreground – the embedding from the standard average self-attention presents improved performance. On all other benchmarks, our approach proves to be the most effective, further validating the robustness of our selection method.
@@
Mask Refinement V21 Object
without background cleaning 59.9 37.1
with background cleaning 61.5 41.0
without background cleaning 63.9 40.3
with background cleaning 65.8 45.1
Table 4: Ablation study on the impact of the background cleaning procedure. We report the results using DINOv2 ViT-B.
Effect of Background Cleaning. Table 4 shows how the performance is affected by the background cleaning mechanism and by the usage of PAMR for mask refinement. It can be observed that the background cleaning procedure has a significantly positive impact on Pascal VOC and COCO Object, leading, respectively, to a +1.6 and +3.9 increase in mIoU score. Further, it can be noticed that the effectiveness of the proposed background cleaning procedure is confirmed also when applying the mask refinement. Qualitative results showing the effect of the background cleaning procedure are available in the supplementary material.
Qualitative Results. Fig. 4 depicts qualitative segmentation results, in which we highlight the segmentation capabilities of Talk2DINO along with other state-of-the-art models (i.e., FreeDA [3], ProxyCLIP [26], and CLIP-DINOiser [52]). We show two images from Pascal VOC, in which it can also be appreciated how the background cleaning procedure leads to high-quality masks and localization, and two images from COCO Stuff and Pascal Context where Talk2DINO effectively segments “things” in the scene, such as the boat and teddy bear, and “stuff” categories, such as sky and road.
@@ 5 Conclusion
In this paper, we introduced Talk2DINO, a novel approach for OVS that bridges the spatially detailed embeddings of the DINOv2 self-supervised vision backbone with the highly semantic text embeddings of CLIP. Our method achieves fine-grained alignment between textual concepts and visual patches without the need for extensive fine-tuning of the backbone networks, only leveraging self-attention maps from DINOv2 and a lightweight language-to-vision mapping layer. Talk2DINO achieves state-of-the-art results in standard OVS benchmarks, outperforming previous prototype-based, CLIP adaptation, and contrastive learning approaches. Our study highlights the potential of integrating vision-only and multimodal models, suggesting broader applications in fine-grained and cross-modal tasks.
@@ Acknowledgments
Figure 5: Performance vs. Parameter Count. The y-axis denotes the obtained mIoU averaged over all the five benchmarks reported in the main paper. Dashed lines connect methods tested with multiple backbone sizes. These are labeled S/B/L for Small/Base/Large ViT models. Talk2DINO offers the best trade-off between performance and number of parameters.
In the following, we present additional material and experimental analyses of the proposed Talk2DINO approach.
@@ Appendix A Additional Experiments and Analyses
Analysis of Model Parameters. Fig. 5 reports a comparison of the relationship between the performance, in terms of average mIoU, and the number of parameters of the models. As it can be observed, Talk2DINO presents a lower number of parameters than the recent competitors FreeDA [3] and ProxyCLIP [26], along with an improved average mIoU. Models with a comparable number of parameters, such as TCL [9], GroupViT [54], and MaskCLIP [63], exhibit a lower performance compared to Talk2DINO. Finally, it shall be noted that models such as FreeDA and ReCo [41] require maintaining external sources of knowledge, which increases memory consumption. Further discussion on the comparison between Talk2DINO, ProxyCLIP, and FreeDA can be found in the following sections (see ”Comparison with ProxyCLIP and FreeDA”).
Role of DINO Registers. The main configuration of Talk2DINO, with both the base and large sizes, leverages the variant of DINOv2 with registers. In Fig. 8 we depict, on the first row, the average self-attentions between the CLS and the other tokens for the ViT-S, ViT-B, and ViT-L architectures with and without registers, while in the following rows, we show the various self-attention heads for each backbone. It can be observed that in the ViT-S the artifacts are not present, and the average self-attention between the model with and without the registers is nearly identical. Instead, the ViT-B exhibits artifacts in the top left corner, resulting in an average self-attention that is especially focused on that portion of the image. This side effect is even more noticeable with the ViT-L, for which the artifact is the only visible token in the average self-attention. These observations align with the results reported in Tab. 2, that show a downgrade in performance without the registers that is directly related to the presence of the artifacts in the self-attentions. Indeed, the largest difference in performance is measured in the ViT-L architecture, while in the ViT-S case, the backbone without registers performs better on four benchmarks out of five.
@@
Choosing Different Visual backbones. In Table 2 of the main paper, we report the performance of our approach applied to different visual backbones. The results demonstrated that our training pipeline is especially suitable for DINOv2, while it leads to unsatisfactory performance on DINO, MAE, and CLIP. We attribute this performance gap to two major factors: (i) the quality of the attention maps and (ii) the semantic richness of the patch representations.
For the first point, we qualitatively analyze the self-attention patterns of the different backbones. Fig. 9 showcases the average self-attentions between the CLS token and the other tokens in the first row, breaking down the contributions from the various self-attention heads in the successive rows. We observe that the self-attention heads of CLIP introduce a noise pattern similar to what we observed for DINOv2 without registers, which limits the effectiveness of our training pipeline. On the other hand, the self-attention maps of DINO and MAE appear cleaner and emphasize homogeneous image regions. However, in these cases, the performance gap with DINOv2 can be attributed to the insufficient semantic richness of the extracted dense features.
To quantitatively assess the patch-level semantics of these backbones, we conduct an experiment in which we classify each patch through linear probing on the images of VOC. We determine the ground-truth labels of the patches via majority voting and evaluate accuracy on the validation set (batch size = 16, learning rate = $5\times 10^{-3}$, for 3 epochs, with 32×32 patches per image, using ViT-B as the backbone). The results, reported in Table 5, align with the overall trends highlighted in the paper: DINOv2 consistently emerges as the best-performing backbone, MAE as the worst, and CLIP and DINO as intermediate. These findings further confirm that the semantic richness of the features extracted by different backbones plays a crucial role in the effectiveness of our approach. Similar conclusions were drawn in the ablation study of FreeDA [3], where a comparable performance drop was observed when using CLIP or DINO instead of DINOv2.
Visual Backbone ViT-S ViT-B ViT-L
MAE - 0.56 0.56
CLIP - 0.89 0.89
DINO 0.62 0.73 -
DINOv2 (without registers) 0.95 0.96 0.95
DINOv2 (with registers) 0.96 0.97 0.96
Table 5: Patch linear probing accuracy on VOC for ViT-Small, ViT-Base, and ViT-Large.
@@
V20 C59 Stuff City ADE
Effect of Training CLIP Last Layer
Trained 77.9 31.5 21.3 34.6 18.7
Frozen 87.1 39.8 28.1 36.6 21.1
Effect of Text Token Selection
Average text tokens 84.7 37.9 25.7 33.6 20.0
Text token to best self-attn map 83.9 33.8 24.2 29.5 18.1
Text token to best self-attn map (NS) 80.8 33.9 23.7 27.5 18.6
CLS token only 87.1 39.8 28.1 36.6 21.1